OpenStudy (anonymous):

If f(x)=(x-1)(x^2+2)^3, then f'(x)=? I did this so far: u=x-1, u'=1, v=(x^2+2)^3, v'=6x(x^2+2)^2 Now when I use the product rule, I keep getting the wrong answer.

OpenStudy (anonymous):

it should work, show me how you worked out the product rule

OpenStudy (anonymous):

(x-1)(6x(x^2+1)^2)+(x^2+2)^3 Now I don't know which one to expand or multiply If i open parentheses, I get (x-1)(x^4+2x^2+1)+(x^6+6x^4+12x^2+8) =x^5+2x^3+x-x^4-2x^2-1+x^6+6x^4-12x^2+8 =x^6+x^5+5x^4+2x^3-24x^2+x+7 the choices are A) 6x(x^2+2)^2 B) 6x(x-1)(x^2+2)^2 C) (x^2+2)^2 (x^2+3x-1) D) (x^2+2)^2 (7x^2-6x+2) E) -3(x-1)(x^2+2)^2

OpenStudy (anonymous):

it is d because \[\large (x-1)*6x*(x^2+2)^2+(x^2+2)^3=(x^2+2)^2 (7x^2-6x+2)\]

OpenStudy (anonymous):

but it's pretty long to write it all down...

OpenStudy (anonymous):

so which parentheses did you expand? how did you get 7x^2−6x+2?

OpenStudy (anonymous):

first factor out (x^2+2)^2 from both sides

OpenStudy (anonymous):

so you'll be left with \( \large (x^2+2)^2((x-1)6x+(x^2+2)^1) \)

OpenStudy (anonymous):

then distribute the 6x over x-1

OpenStudy (anonymous):

\( \large (x^2+2)^2(6x^2-6x+x^2+2) \)

OpenStudy (anonymous):

Okay I got it now. Thank you!

OpenStudy (anonymous):

no problem